OpenStudy (anonymous):

how to solve 3(x+4)=-12

OpenStudy (anonymous):

x+4=-4 x=-8

OpenStudy (anonymous):

3x + 12 = -12 3x = -24 x = -8

OpenStudy (anonymous):

3x+12= -12 3x= -12-12 3x= -24 x= -24/3 = -8